OpenStudy (paradise):

Solve the following system of equations by using the elimination method. x – y = 11 2x + y = 19

OpenStudy (paradise):

@nirmalnema here it is

OpenStudy (nirmalnema):

subtract 1 eq from the 2 nd one

OpenStudy (nirmalnema):

then y will be eliminated and you will get the expression in terms of x

OpenStudy (nirmalnema):

solve it to get x..and then put the value of x is any one eq u get the value of y

OpenStudy (nirmalnema):

here is the solution x-y=11 2x+y=19 so after subtraction 3x=30 and x = 10 therefore 10-y=11 y = -1

OpenStudy (nirmalnema):

@Paradise sorry for replying so late i was been blocked thats y i was unable to ans it...i m really sorry